# Hybrid Quantum Repeater Chains with Semiconductor Quantum Dots and Group-IV-Vacancy Color Centers in Diamond

### v2

@@ III System compatibility
We operate in the weak-driving regime, where the product of the photon’s electric field amplitude and the G4V transition dipole moments is much weaker than the spin’s optical decay rates $\{\gamma_{1A},\gamma_{2B}\}$. In this regime, the photon induces no population transfer but imparts a conditional phase on the spin qubit, implementing a photon-induced phase gate [32].
@@
Since spectrally degenerate photons are resonant with only one of the optical transitions, simply applying identical phase gates to both time bins would not generate entanglement. However, entanglement can still be created by inserting a spin rotation between the two photon arrivals [see Fig. 3(e)], provided that the spin rotation does not commute with the phase gates [60, 54]. Here, we consider a $\pi/2$ rotation around the $y$-axis acting on the G4V spin, which can be implemented via a microwave pulse [61] with an error rate below $10^{-4}$ [62]. Following a projective measurement of the photonic qubit, the photonic state is effectively transferred to the spin. Details are provided in App. A.
@@
When evaluating the objective function in Eq. (2), we account for the broadband nature of the QD photons. A detailed analysis of the associated effects is presented in App. B. To enhance the frequency contrast between the $|1\rangle\leftrightarrow|A\rangle$ and $|2\rangle\leftrightarrow|B\rangle$ transitions, we apply an off-axis 1 T magnetic field to both the SiV and SnV centers, which we investigate as light and heavy representatives of G4V, respectively. In addition, a compressive strain of $E_{x}=5.38\cdot 10^{-5}$, perpendicular to the G4V symmetry axis, is applied to the SiV but not the SnV as an outcome of parameter optimization [61] (see App. D.1 for details). As the angle between the magnetic field and the symmetry axis increases, the spectral contrast between the intended transitions improves, leading to better performance. However, this also increases cross-coupling, the unintended coupling between transitions $|1\rangle\leftrightarrow|B\rangle,|2\rangle\leftrightarrow|A\rangle$, which introduces nonlinearities in the system’s equations of motion. These are included in our modeling and additionally utilized to minimize entanglement infidelities (see App. D.2). Such cross-coupling terms that have not been addressed yet for G4Vs.
@@
We determine optimal magnetic field orientations for both SiV and SnV subject to the given conditions and utilize crosstalk to minimize the spin-spin entanglement infidelity using state-of-the-art optimization techniques for the rates $\gamma_{\rm X}=4.34$ GHz and $\gamma_{\rm XX}=8.33$ GHz (see App. D.2). Assuming a perfect photonic Bell state and no filtering, the resulting minimum infidelity and success probability for the SiV are $1-F_{\rm sp,SiV}=7.91\cdot 10^{-2}$ and $\eta_{\rm sp,SiV}=0.9187$, achieved by a SiV-cooperativity $C_{\rm SiV}=\left.g_{1A}^{2}/(2\gamma_{1A}\,\kappa)\right|_{\rm SiV}\approx 11$ (see App. D.2 for the optimized parameters). However, the spin splitting for the considered magnetic field configuration is $\omega_{s}\approx 26$ GHz, which is challenging to address using microwave spin control. To address realistic parameters, we also consider a magnetic field strength that produces a spin splitting closer to the experimentally achieved value [62]. We choose $B=0.3$ T yielding $\omega_{s}\approx 8$ GHz. One limitation of the reduced magnetic field strength is a reduction of the maximum feasible bandwidth of the QD photons.
@@
For the SnV, we obtain $1-F_{\rm sp,SnV}=6.09\cdot 10^{-2}$, $\eta_{\rm sp,SnV}=0.9935$, and $C_{\rm SnV}=160$. A higher cooperativity $C_{\rm SnV}$ compared to $C_{\rm SiV}$ is required due to the much smaller natural atomic decay rate $\gamma_{1A,\mathrm{SnV}}$. Since such a high cooperativity for the SnV has not yet been experimentally demonstrated, we evaluate the case where $C_{\rm SnV}=25$, matching already experimentally demonstrated values [54]. For this relatively low cooperativity, a bandwidth as low as $\gamma<480$ MHz is required to achieve an infidelity similar to the SiV’s spin-spin entangled state infidelity $1-F_{\rm sp}<7.9\cdot 10^{-2}$. We conclude that broadband photons, with rates comparable to those of the quantum dot considered here, are compatible with future SnV cavities with cooperativities on the order of $100$. Achieving such high cooperativities may be enabled by advances in high-purity diamond growth [63], isotopic purification [64], and optimized surface termination [65]. Current state-of-the-art SnV cavities can already interface with broadband QDs through bandwidth compression [66] or spectral filtering. Alternatively, one could employ a QD with intrinsically narrower bandwidth.
Next, we examine the infidelity $1-F_{\rm sp}$ and efficiency $\eta_{\rm sp}$ as a function of $\gamma_{XX}$ fixing $\gamma_{X}=4.34$ GHz for a given magnetic and strain field induced optical splitting using frequency filtering, which we show in Fig. 4. The infidelity increases slightly with bandwidth and then rises sharply at larger bandwidths. This behavior reflects the finite contrast of the system; since the cavity and frequency of the incoming photons were previously optimized based on the QD properties, the steep increase occurs near that range. Below we go more into detail of that behavior. As the bandwidth increases, the phase contrast between $|1\rangle\leftrightarrow|A\rangle$ and $|2\rangle\leftrightarrow|B\rangle$ is washed out as was explained in [33]. The phase contrast can be recovered by increasing the optical splitting $\Delta\omega_{s}$, which depends on the magnetic field strength, orientation, and strain configuration. Interestingly, the cooperativity cannot be increased indefinitely at a given splitting for increasing the infidelity. This arises because the phase contrast results from a competition between the contrast itself and its first-order derivative with respect to the cooperativity [33]. As higher cooperativities reduce the magnitude of the first-order derivative, the overall fidelity increases at the price of a reduced phase difference.
@@
Lastly, we analyze the spin-spin entanglement fidelity at the optimized cavity and magnetic field configuration in a local environment of the optimized cavity parameters. For a deviation of maximally 11 GHz for the cavity loss rate $\kappa$ and cavity mode detuning $\delta:=\omega_{1A}-\omega_{c}$ the spin-spin entanglement infidelity for the SiV (SnV) stays below $8.11\cdot 10^{-2}$ $(10.0\cdot 10^{-2})$ and efficiency above $0.9175$ $(0.9931)$ (see App. D.3 for details). For this particular example, we find that the SiV’s spin–spin entangled state is more robust to fabrication uncertainties than the SnV’s. This does not need to be generally true, but only for this set of local minima of the optimization. We conclude that the spin-spin entanglement fidelity at the optimized cavity and magnetic field configuration is also robust to fabrication uncertainties.
@@ IV Repeater chain Protocol
Figure 5: Operational protocol and optimal secret-key rate for a QD-SiV quantum repeater chain. a) Operation timeline of the repeater chain. The red, blue, and yellow boxes represent the qubit-loading phase, one round of entanglement-distillation operation, and the entanglement swapping phase. Each chain of boxes represents an operational thread of a memory cell. b) Maximal secret-key rate ($R_{\rm sk,opt}$) as a function of total distance $L$ under certain memory cell number per module ($m$), QD photon bandwidth presented as spin Bell pair fidelity ($F_{\rm sp}$), and the photon-mediated nuclear-nuclear gate error rate $\varepsilon_{\rm n-n}$. The spin Bell pair fidelities $F_{\rm sp}=0.95$ and $0.98$ correspond to the QD photon bandwidth after the frequency filtering of $\tilde{\gamma}_{\rm X},\tilde{\gamma}_{\rm XX}]=[4.34,8.17]$ GHz and $\tilde{\gamma}_{\rm X},\tilde{\gamma}_{\rm XX}]=[4.33,6.50]~\text{GHz}$, respectively. Note that for the $F_{\rm sp}=0.95$ case, only one photon in a Bell pair is filtered ($\tilde{\gamma}_{\rm X}=\gamma_{\rm X}=4.34~\text{GHz}$). The optimal values of NN, $n_{\rm loa}$, $n_{\rm dis,n}$, and $n_{\rm dis,e}$, which give the maximal secret-key rate, can be found in App. F.
@@
QD emitter duty cycle ($t_{\rm QD}$) 1 ns [49, 25, 50]
nuclear spin coherence time ($T_{nu}$) 0.1 s [67, 68, 69]
fiber attenuation rate ($\gamma_{\rm fib}$) 0.2 dB/km [18]
@@
G4V electron spin reset time ($t_{\rm res}$) 1 $\mu$[31]
electron-nuclear gate time ($t_{\rm nu}$) 10 $\mu$[67, 68, 69]
@@
$5\times 10^{-5}$ [67, 68, 69]
@@
cavity-fiber coupling efficiency ($\eta_{\rm c-f}$) 0.864 [54]
@@
0.98 [31]
frequency converter efficiency ($\eta_{\rm fc}$) 0.73 [70]
photon detection efficiency ($\eta_{\rm pd}$) 0.99 [54]
@@
0.83 [71]
optical switch efficiency ($\eta_{\rm swi}$) 0.95 [72]
To assess the repeater chain performance, we consider a symmetric node configuration and a cyclical operational protocol. This configuration and protocol may not be optimal, but they provide a lower bound on the secret-key rate. Specifically, we consider N1N-1 repeater nodes that evenly divide the total communication distance $L$ into NN segments. Each repeater node contains two quantum memory modules, used to establish elementary entanglement links with adjacent segments (see Fig. 1). The two end nodes have either one or zero memory modules, depending on whether entanglement distillation is applied. When no entanglement distillation is applied, the incoming photons are measured directly, so no quantum memories are required. Each memory module consists of $m$ memory cells, with each cell composed of a G4V electron spin and a nearby nuclear spin. The nuclear spin serves as long-term storage for a single qubit, while the electron spin provides the interface to photonic qubits. Here, we consider $m\in\{100,1{,}000\}$.
@@
where $t_{\rm QD}=1\,\text{ns}$ [49, 25, 50] is the duty cycle of the QD emitter.
@@
where $c=2\times 10^{5}$ km/s is the transmission speed of a classical signal. At the same time, the qubit is transferred from the electron spin to the nuclear spin, which has a longer coherence time $T_{\rm nu}=0.1\,\text{s}$ [67, 68, 69]. The electron-nuclear transfer requires $t_{\rm nu}=10\,\mu\text{s}$ [67, 68, 69]. Therefore, the duration of the qubit-loading phase is
@@
is the probability of loading one qubit into one arm of a segment, where $\gamma_{\rm fib}=0.2$ dB/km [18] is the fiber attenuation rate, and
@@
As mentioned above, we adopt the entanglement distillation protocol of Ref. [1]. In this protocol, one can distill raw Bell pairs to the first, second, or third level. Level-1, -2, and -3 distillation consume two, four, and six raw Bell pairs, respectively, and require one, two, and three sequential distillation rounds, as shown in Fig. 1c. Each round involves similar operations: single-qubit rotations on the nuclear spins, intra-module nucleus–nucleus gates, nuclear-spin readout, and cross-module classical communication.
Similar to the entanglement-swapping procedure described in Sec. II.2, we employ two G4V electron spins as the medium to implement nucleus–nucleus gates. The procedure is as follows: first, establish entanglement between the two electron spins; then perform local electron–nucleus gates. The electron–electron entanglement is photon-mediated and attempted $n_{\rm e-e}$ times. Each attempt consumes time $t_{\rm res}$ and carries a probability $\varepsilon_{\rm nu}=5\times 10^{-5}$ [67, 68, 69] of inducing nuclear-spin decoherence. The success probability per attempt is
@@
where $|1/2\rangle$ and $|+/$-$\rangle$ are the $Z$- and $X$-basis states of the end-node qubit. According to the BB84 protocol [34], the secret-key rate per entangled pair is
@@
where $H(\cdot)$ is the binary entropy [73]. The ultimate secret-key rate is therefore

### v3

@@ III System compatibility
We operate in the weak-driving regime, where the product of the photon’s electric field amplitude and the G4V transition dipole moments is much weaker than the spin’s optical decay rates $\{\gamma_{1A},\gamma_{2B}\}$. In this regime, the photon induces no population transfer but imparts a conditional phase on the spin qubit, implementing a photon-induced phase gate [strocka_memory_2025].
@@
Since spectrally degenerate photons are resonant with only one of the optical transitions, simply applying identical phase gates to both time bins would not generate entanglement. However, entanglement can still be created by inserting a spin rotation between the two photon arrivals [see Fig. 3(e)], provided that the spin rotation does not commute with the phase gates [borregaard_one-way_2020, bhaskar_experimental_2020]. Here, we consider a $\pi/2$ rotation around the $y$-axis acting on the G4V spin, which can be implemented via a microwave pulse [pieplow_efficient_2024] with an error rate below $10^{-4}$ [karapatzakis_microwave_2024]. Following a projective measurement of the photonic qubit, the photonic state is effectively transferred to the spin. Details are provided in App. A.
@@
When evaluating the objective function in Eq. (2), we account for the broadband nature of the QD photons. A detailed analysis of the associated effects is presented in App. B. To enhance the frequency contrast between the $|1\rangle\leftrightarrow|A\rangle$ and $|2\rangle\leftrightarrow|B\rangle$ transitions, we apply an off-axis 1 T magnetic field to both the SiV and SnV centers, which we investigate as light and heavy representatives of G4V, respectively. In addition, a compressive strain of $E_{x}=5.38\cdot 10^{-5}$, perpendicular to the G4V symmetry axis, is applied to the SiV but not the SnV as an outcome of parameter optimization [pieplow_efficient_2024] (see App. D.1 for details). As the angle between the magnetic field and the symmetry axis increases, the spectral contrast between the intended transitions improves, leading to better performance. However, this also increases cross-coupling, the unintended coupling between transitions $|1\rangle\leftrightarrow|B\rangle,|2\rangle\leftrightarrow|A\rangle$, which introduces nonlinearities in the system’s equations of motion. These are included in our modeling and additionally utilized to minimize entanglement infidelities (see App. D.2). Such cross-coupling terms that have not been addressed yet for G4Vs.
@@
We determine optimal magnetic field orientations for both SiV and SnV subject to the given conditions and utilize crosstalk to minimize the spin-spin entanglement infidelity using state-of-the-art optimization techniques for the rates $\gamma_{\rm X}=4.34$ GHz and $\gamma_{\rm XX}=8.33$ GHz (see App. D.2). Assuming a perfect photonic Bell state and no filtering, the resulting minimum infidelity and success probability for the SiV are $1-F_{\rm sp,SiV}=7.91\cdot 10^{-2}$ and $\eta_{\rm sp,SiV}=0.9187$, achieved by a SiV-cooperativity $C_{\rm SiV}=\left.g_{1A}^{2}/(2\gamma_{1A}\,\kappa)\right|_{\rm SiV}\approx 11$ (see App. D.2 for the optimized parameters). However, the spin splitting for the considered magnetic field configuration is $\omega_{s}\approx 26$ GHz, which is challenging to address using microwave spin control. To address realistic parameters, we also consider a magnetic field strength that produces a spin splitting closer to the experimentally achieved value [karapatzakis_microwave_2024]. We choose $B=0.3$ T yielding $\omega_{s}\approx 8$ GHz. One limitation of the reduced magnetic field strength is a reduction of the maximum feasible bandwidth of the QD photons.
@@
For the SnV, we obtain $1-F_{\rm sp,SnV}=6.09\cdot 10^{-2}$, $\eta_{\rm sp,SnV}=0.9935$, and $C_{\rm SnV}=160$. A higher cooperativity $C_{\rm SnV}$ compared to $C_{\rm SiV}$ is required due to the much smaller natural atomic decay rate $\gamma_{1A,\mathrm{SnV}}$. Since such a high cooperativity for the SnV has not yet been experimentally demonstrated, we evaluate the case where $C_{\rm SnV}=25$, matching already experimentally demonstrated values [bhaskar_experimental_2020]. For this relatively low cooperativity, a bandwidth as low as $\gamma<480$ MHz is required to achieve an infidelity similar to the SiV’s spin-spin entangled state infidelity $1-F_{\rm sp}<7.9\cdot 10^{-2}$. We conclude that broadband photons, with rates comparable to those of the quantum dot considered here, are compatible with future SnV cavities with cooperativities on the order of $100$. Achieving such high cooperativities may be enabled by advances in high-purity diamond growth [Chakraborty2019], isotopic purification [Teraji2015], and optimized surface termination [Hauf2011]. Current state-of-the-art SnV cavities can already interface with broadband QDs through bandwidth compression [Allgaier2017] or spectral filtering. Alternatively, one could employ a QD with intrinsically narrower bandwidth.
Next, we examine the infidelity $1-F_{\rm sp}$ and efficiency $\eta_{\rm sp}$ as a function of $\gamma_{XX}$ fixing $\gamma_{X}=4.34$ GHz for a given magnetic and strain field induced optical splitting using frequency filtering, which we show in Fig. 4. The infidelity increases slightly with bandwidth and then rises sharply at larger bandwidths. This behavior reflects the finite contrast of the system; since the cavity and frequency of the incoming photons were previously optimized based on the QD properties, the steep increase occurs near that range. Below we go more into detail of that behavior. As the bandwidth increases, the phase contrast between $|1\rangle\leftrightarrow|A\rangle$ and $|2\rangle\leftrightarrow|B\rangle$ is washed out as was explained in [omlor_entanglement_2024]. The phase contrast can be recovered by increasing the optical splitting $\Delta\omega_{s}$, which depends on the magnetic field strength, orientation, and strain configuration. Interestingly, the cooperativity cannot be increased indefinitely at a given splitting for increasing the infidelity. This arises because the phase contrast results from a competition between the contrast itself and its first-order derivative with respect to the cooperativity [omlor_entanglement_2024]. As higher cooperativities reduce the magnitude of the first-order derivative, the overall fidelity increases at the price of a reduced phase difference.
@@
Lastly, we analyze the spin-spin entanglement fidelity at the optimized cavity and magnetic field configuration in a local environment of the optimized cavity parameters. For a deviation of maximally $1$ GHz for the cavity loss rate $\kappa$ and cavity mode detuning $\delta:=\omega_{1A}-\omega_{c}$ the spin-spin entanglement infidelity for the SiV (SnV) stays below $8.11\cdot 10^{-2}$ $(10.0\cdot 10^{-2})$ and efficiency above $0.9175$ $(0.9931)$ (see App. D.3 for details). For this particular example, we find that the SiV’s spin–spin entangled state is more robust to fabrication uncertainties than the SnV’s. This does not need to be generally true, but only for this set of local minima of the optimization. We conclude that the spin-spin entanglement fidelity at the optimized cavity and magnetic field configuration is also robust to fabrication uncertainties.
@@ IV Repeater chain Protocol
Figure 5: Operational protocol and optimal secret-key rate for a QD-SiV quantum repeater chain. a) Operation timeline of the repeater chain. The red, blue, and yellow boxes represent the qubit-loading phase, one round of entanglement-distillation operation, and the entanglement swapping phase. Each chain of boxes represents an operational thread of a memory cell. b) Maximal secret-key rate ($R_{\rm sk,opt}$) as a function of total distance $L$ under certain memory cell number per module ($m$), QD photon bandwidth presented as spin Bell pair fidelity ($F_{\rm sp}$), and the photon-mediated nuclear-nuclear gate error rate $\varepsilon_{\rm n-n}$. The spin Bell pair fidelities $F_{\rm sp}=0.95$ and $0.98$ correspond to the QD photon bandwidth after the frequency filtering of $\tilde{\gamma}_{\rm X},\tilde{\gamma}_{\rm XX}]=[4.34,8.17]$ GHz and $\tilde{\gamma}_{\rm X},\tilde{\gamma}_{\rm XX}]=[4.33,6.50]~\text{GHz}$, respectively. Note that for the $F_{\rm sp}=0.95$ case, only one photon in a Bell pair is filtered ($\tilde{\gamma}_{\rm X}=\gamma_{\rm X}=4.34~\text{GHz}$). The optimal values of $N$, $n_{\rm loa}$, $n_{\rm dis,n}$, and $n_{\rm dis,e}$, which give the maximal secret-key rate, can be found in App. F.
@@
QD emitter duty cycle ($t_{\rm QD}$) 1 ns [QD_4_strobel2024high, schimpf_quantum_2021, Huber2018]
nuclear spin coherence time ($T_{nu}$) 0.1 s [SiV_readout1_knaut2024entanglement, SiV_readout2_wei2024universal, SiV_readout3_stas2022robust]
fiber attenuation rate ($\gamma_{\rm fib}$) 0.2 dB/km [miya_ultimate_1979]
@@
G4V electron spin reset time ($t_{\rm res}$) 1 $\mu$[bopp_sawfish_2024]
electron-nuclear gate time ($t_{\rm nu}$) 10 $\mu$[SiV_readout1_knaut2024entanglement, SiV_readout2_wei2024universal, SiV_readout3_stas2022robust]
@@
$5\times 10^{-5}$ [SiV_readout1_knaut2024entanglement, SiV_readout2_wei2024universal, SiV_readout3_stas2022robust]
@@
cavity-fiber coupling efficiency ($\eta_{\rm c-f}$) 0.864 [bhaskar_experimental_2020]
@@
0.98 [bopp_sawfish_2024]
frequency converter efficiency ($\eta_{\rm fc}$) 0.73 [Zaske2011]
photon detection efficiency ($\eta_{\rm pd}$) 0.99 [bhaskar_experimental_2020]
@@
0.83 [thorlabs]
optical switch efficiency ($\eta_{\rm swi}$) 0.95 [Zeng2024]
To assess the repeater chain performance, we consider a symmetric node configuration and a cyclical operational protocol. This configuration and protocol may not be optimal, but they provide a lower bound on the secret-key rate. Specifically, we consider N1N-1 repeater nodes that evenly divide the total communication distance $L$ into $N$ segments. Each repeater node contains two quantum memory modules, used to establish elementary entanglement links with adjacent segments (see Fig. 1). The two end nodes have either one or zero memory modules, depending on whether entanglement distillation is applied. When no entanglement distillation is applied, the incoming photons are measured directly, so no quantum memories are required. Each memory module consists of $m$ memory cells, with each cell composed of a G4V electron spin and a nearby nuclear spin. The nuclear spin serves as long-term storage for a single qubit, while the electron spin provides the interface to photonic qubits. Here, we consider $m\in\{100,1{,}000\}$.
@@
where $t_{\rm QD}=1\,\text{ns}$ [QD_4_strobel2024high, schimpf_quantum_2021, Huber2018] is the duty cycle of the QD emitter.
@@
where $c=2\times 10^{5}$ km/s is the transmission speed of a classical signal. At the same time, the qubit is transferred from the electron spin to the nuclear spin, which has a longer coherence time $T_{\rm nu}=0.1\,\text{s}$ [SiV_readout1_knaut2024entanglement, SiV_readout2_wei2024universal, SiV_readout3_stas2022robust]. The electron-nuclear transfer requires $t_{\rm nu}=10\,\mu\text{s}$ [SiV_readout1_knaut2024entanglement, SiV_readout2_wei2024universal, SiV_readout3_stas2022robust]. Therefore, the duration of the qubit-loading phase is
@@
is the probability of loading one qubit into one arm of a segment, where $\gamma_{\rm fib}=0.2$ dB/km [miya_ultimate_1979] is the fiber attenuation rate, and
@@
As mentioned above, we adopt the entanglement distillation protocol of Ref. [distillation_nigmatullin2016minimally]. In this protocol, one can distill raw Bell pairs to the first, second, or third level. Level-1, -2, and -3 distillation consume two, four, and six raw Bell pairs, respectively, and require one, two, and three sequential distillation rounds, as shown in Fig. 1c. Each round involves similar operations: single-qubit rotations on the nuclear spins, intra-module nucleus–nucleus gates, nuclear-spin readout, and cross-module classical communication.
Similar to the entanglement-swapping procedure described in Sec. II.2, we employ two G4V electron spins as the medium to implement nucleus–nucleus gates. The procedure is as follows: first, establish entanglement between the two electron spins; then perform local electron–nucleus gates. The electron–electron entanglement is photon-mediated and attempted $n_{\rm e-e}$ times. Each attempt consumes time $t_{\rm res}$ and carries a probability $\varepsilon_{\rm nu}=5\times 10^{-5}$ [SiV_readout1_knaut2024entanglement, SiV_readout2_wei2024universal, SiV_readout3_stas2022robust] of inducing nuclear-spin decoherence. The success probability per attempt is
@@
where $|1/2\rangle$ and $|+/$-$\rangle$ are the $Z$- and $X$-basis states of the end-node qubit. According to the BB84 protocol [shor_simple_2000], the secret-key rate per entangled pair is
@@
where $H(\cdot)$ is the binary entropy [noauthor_david_2016]. The ultimate secret-key rate is therefore